### v1

@@ Abstract
Object-Centric Process Mining (OCPM) enables business process analysis from multiple perspectives. For example, an educational path can be examined from the viewpoints of students, teachers, and groups. This analysis depends on Object-Centric Event Data (OCED), which captures relationships between events and object types, representing different perspectives. Unlike traditional process mining techniques, extracting OCED minimizes the need for repeated log extractions when shifting the analytical focus. However, recording these complex relationships increases the complexity of the log extraction process. To address this challenge, this paper proposes a method for extracting OCED based on PM, a well-established process mining framework. Our approach introduces a structured framework that guides data analysts and engineers in extracting OCED for process analysis. We validate this framework by applying it in a real-world educational setting, demonstrating its effectiveness in extracting an Object-Centric Event Log (OCEL), which serves as the standard format for recording OCED, from a learning management system and an administrative grading system.
@@ 2.1 Process Mining Methodologies
Process Diagnostics Method (PDM) [4], the L* Life-Cycle Model [5], and PM2 [18] are three well-known process mining methodologies followed by different people to apply process mining in practice.
The Process Diagnostics Method (PDM) [4] focuses on rapid, high-level analysis of business processes without requiring domain-specific knowledge. It examines control flow, performance, and organizational aspects based on event logs, providing quick insights. However, PDM lacks explicit iterative refinement of the analysis [18], and is less suitable for long-term process optimization efforts [4].
@@ 2.2 Brief Overview of OCED
(a) OCEL 2.0 metamodel
(b) EKG metamodel
Figure 1: Two sample metamodels for storing OCED
The OCEL 2.0 metamodel, shown on the left side of the figure in UML notation, defines an event as an atomic action that occurs in a business process, representing the execution of an event type (a.k.a, an activity), such as “place order” in Order-to-cash process. Each event type can have multiple event attributes, and each event can store different event attribute values. OCEL records objects as separate entities, where each object belongs to an object type. For example, the “order” can be an object type with multiple instances representing different orders. Each object type can have multiple object attributes, and each object can store different values for these attributes at specific points in time. Objects can also be related to other objects through object-to-object relations; for example, an order can contain multiple items. As illustrated in the metamodel, events can be associated with multiple objects and vice versa through event-to-object relations.
@@ 2.3 OCED Extraction Initiatives
Beyond SAP-specific solutions, Xiong et al. introduced a method for extracting OCEL from relational databases using a Virtual Knowledge Graph (VKG) approach [19]. Their method extended the OnProm framework to support both OCEL and the traditional XES format. By leveraging ontology-based data access (OBDA) and the VKG system Ontop, they defined domain ontologies and mappings, enabling event log extraction via SPARQL queries. Their approach was validated using the Dolibarr ERP system, demonstrating how relational data can be transformed into structured OCEL logs.
@@ 3 The OCPM2 Methodology
This section introduces OCPM2, an Object-Centric Process Mining methodology defined as an extension of PM2 [18] to facilitate the extraction of Object-Centric Event Data (OCED) [7]. We define our method by extending PM2 to ensure the reuse of a well-established and previously tested methodology for event log extraction, originally designed for traditional process mining. Our approach retains the clarity and efficiency of PM2 while making it suitable for extracting object-centric event data from diverse information systems.
The methodology consists of twelve stages, where the related stages are grouped into phases, as indicated by dashed rectangles in Figure 2. The process begins with the Planning stage, followed by the Domain Modeling, Log Extraction, and Analysis Iteration phases, and concludes with the Process Improvement & Support stage. Each stage produces specific artifacts, as outlined in the figure. The following sections elaborate on these stages, the generated artifacts, and the defined phases. The filled-colored phases and stages are identical to those in the PM2 methodology briefly described in Section 2 (we refer readers to the PM2 [18] for further details.)
Figure 2: An overview of the OCPM2 methodology extending the PM2 methodology
@@ 3.2 Domain Modeling
The Domain Modeling phase consists of five stages (stages 2 to 6 in Figure 2), identifying and conceptualizing the information needed for the extraction of OCED.
@@
The phase begins with Object Type Identification to determine all object types needed to address the questions elicited in the Planing. Because an object type may be required for multiple questions, we propose documenting this using a Question-to-Object Type Matrix (Q2OT). This matrix represents questions as rows and object types as columns, marking intersection cells to indicate the required object types for each question.
@@
Next, in the Conceptual Modeling stage, process analysts and business experts collaborate to define a model ( CM) that captures the relationships between object types. Business experts provide domain knowledge, while process analysts structure this information using a conceptual modeling language to represent key relationships. Among these, recognizing the “is-a” relationship enables capturing OLAP operations such as drill-down and roll-up in OCPM based on OCEL [13, 12].
@@
The Activity Identification stage follows, eliciting different activities (or event types) to which the identified object types and questions are related.
@@
The Extraction Matrix Definition stage defines the extraction matrix (X), where the relationships between object types and activities are documented. When an “is-a” relationship exists among object types, separate columns can specify whether an activity relates to the supertype class or a subtype class. Various documentation approaches can be adapted based on specific perspectives. From a system perspective, CRUD (Create, Read, Update, and Delete) operations are suitable for specifying the type of relationship between an activity and an object type. However, from a process perspective, CRUD may not fully capture the dynamic nature of object interactions. Instead, documenting the quantity range of objects for each object type provides a more process-oriented approach. This method captures the expected number of objects per event and supports data validation by ensuring that the extracted log adheres to these expectations. The details of these adaptations can be delegated to design and development teams.
@@
Finally, in the Extraction Prioritization stage, business, process, and system experts collaborate to prioritize the extraction of objects and events based on their i) feasibility, ii) data availability, or iii) importance. Given the potentially large number of object types and activities, extracting all data at once may not be feasible due to data being fragmented across multiple sources and the need to check data quality step-by-step to catch errors early. To manage this, we propose a prioritization artifact as the output of this stage, including a set of prioritized objects and activities ( Pr). Note that new object types can be identified during this phase, requiring the phase to be repeated to have a more holistic understanding of the domain.
@@ 3.3 Log Extraction
The Log Extraction phase consists of three stages (Stages 7 to 9 in Figure 2) as follows:
In the Implementation stage, developers build pipelines to read data from source systems, transform them into the desired OCED format, and load them into destination storage. If OCEL is the target format, objects should be processed first based on their list in the extraction matrix. When an “is-a” relationship exists among object types, we propose storing objects at the supertype level with an attribute distinguishing subtypes. This structure follows the single-table inheritance technique used in Object-Relational Mapping (ORM) [8]. This ensures referential integrity and facilitates efficient querying and aggregation, allowing drill-down operations [12]. Subsequently, developers implement extraction logic for object-to-object relationships, determined by the conceptual model (stage 3), in parallel with extraction logic for events based on the extraction matrix. Once event extraction is complete, they implement extraction logic for event-to-object relationships. Please note that similar mapping can be done for EKG [6] and tEKG [10] as the OCED formats are transformable to each other.
@@
Although the log extraction process has been adapted to generate OCED, the following stages remain identical to PM2 [18] since the subsequent stages of the process mining apply regardless of the log format. Additionally, the extracted OCED can be flattened to enable the use of traditional process discovery techniques.
@@ 3.4 Analysis Iteration
The Analysis Iteration phase in PM2 supports gaining insights from extracted event data through two stages: Mining & Analysis and Evaluation. The Mining & Analysis stage includes process model discovery ( PM) from the input log ( L), conformance checking, and process enhancement based on analytical insights ( A). This can yield compliance findings ( C) and performance findings ( P). The Evaluation stage results in process improvement ideas ( I) or refined/new process-related questions ( ?+), which may prompt further analysis or new data extraction.
@@ 4 Case Study
This section presents the application of the OCPM methodology in executing a process mining project at the Department of Computer and Systems Sciences, Stockholm University. We elaborate on each stage and phase below.
@@ 4.1 Planning
Although we identified numerous questions to guide the project, we present only a subset here to illustrate the application of the methodology. The information systems identified at this stage included 1) the learning management system, Moodle [14], and 2) the website that records students’ grades.
@@ 4.2 Domain Modeling
Conceptual Modeling: Figure 3 presents the conceptual model using UML notation, illustrating the relationships between object types. This model includes additional object types beyond those represented in the Q2OT matrix, reflecting the iterative nature of our approach. For example, the object type Teacher was later identified by reviewing the Moodle documentation as a relevant entity within the process. Additionally, User was introduced as a supertype encompassing both Teacher and Student as explained in Section 2.
@@ 4.3 Log Extraction
Implementation: We set up development and test environments to build and validate pipelines for extracting data from the systems, integrating them based on the extraction matrix, and loading it into an OCEL format. When loading data into OCEL, we defined User as a single object type, with Teacher and Student modeled as subclasses. This approach facilitated the discovery of process models at a general level while enabling drill-down analyses to distinguish teacher- and student-specific process behaviors.
@@
Verification: We developed a script to systematically verify the accuracy of the extracted event data by mapping activities to object types. Figure 4 presents this structured validation process. In this figure, the expected relationship between Group and set assignment grade is missing. This discrepancy arises because Moodle logs a separate set assignment grade event for each student, even when a teacher grades a group assignment collectively. Automated verification helped identify such inconsistencies, highlighting areas for improvement in future log extractions by developers. Furthermore, the relationship between Teacher and Student is not directly visible because the types of objects were recorded at the User level. However, drill-down analysis can reveal these relationships as needed.
@@ 4.4 Analysis Iteration
For Q1, Figure 5 illustrates an Object-Centric Directly-Follows Graph (OC-DFG), discovered by filtering the extracted log, drilling down page objects into individual page names, and unfolding view page events for each page. This graph visualizes the sequential flow of actions during the course from the perspectives of the User and Course object types.
@@
In this process, the page names for given codes are as follow: A1P1: History of BPM, A1P2: Business Process Models, A1P3: Business Process Enactment, A1P4: Syntax and Semantics, A1P5: Control-flow Patterns, A1P6: Business Process Complexity, and L1P1: Introduction to Process Tree. The model reveals that students sequentially accessed five pages: A1P1 to A1P5. Additionally, a loop between A1P5, A1P6, and L1P1 suggests students frequently revisit these pages, indicating potential room for improvement.
@@
Although, due to space constraints, we cannot present all analysis results in detail, extracting OCELs enabled us to investigate all the questions raised while interviewing teachers at the beginning of the project.
@@ 5 Conclusion
To validate our approach, we applied OCPM2 in an educational setting, extracting and analyzing OCELs from a learning management system and an administrative grading system. Our results demonstrated the practical usefulness of OCPM2, i.e., its ability to effectively reveal student navigation patterns and assignment submission processes. The case study also exposed limitations in existing system logs, reinforcing the need for improved object-to-activity relationships in event data.
The case study exposed limitations in system logs, particularly in capturing accurate object-to-activity relationships. A key issue was the missing link between Group and set assignment grade, causing Moodle to log separate grading events per student, even for collectively graded assignments. This discrepancy resulted in a mismatch between the number of students who submitted the assignment and those involved in the grading process. Automated verification identified these inconsistencies, helping developers to extract correct event logs in future extractions.
Future research will focus on enhancing automated verification mechanisms, refining multi-dimensional process analysis operations, and expanding OCPM2’s applicability across other domains, such as healthcare and financial services. OCPM2 contributes to the advancement of Object-Centric Process Mining by providing an orchestrated approach to the mining task, with the potential for further enhancements such as automation and reuse.

### v2

@@ Abstract
Object-Centric Process Mining (OCPM) enables business process analysis from multiple perspectives. For example, an educational path can be examined from the viewpoints of students, teachers, and groups. This analysis depends on Object-Centric Event Data (OCED), which captures relationships between events and object types, representing different perspectives. Unlike traditional process mining techniques, extracting OCED minimizes the need for repeated log extractions when shifting the analytical focus. However, recording these complex relationships increases the complexity of the log extraction process. To address this challenge, this paper proposes a methodology for extracting OCED based on PM, a well-established process mining framework. Our approach introduces a structured framework that guides data analysts and engineers in extracting OCED for process analysis. We validate this framework by applying it in a real-world educational setting, demonstrating its effectiveness in extracting an Object-Centric Event Log (OCEL), which serves as the standard format for recording OCED, from a learning management system and an administrative grading system.
@@ 2.1 Process Mining Methodologies
Process Diagnostics Method (PDM) [4], the L* Life-Cycle Model [5], and PM2 [23] are three well-known process mining methodologies followed by different people to apply process mining in practice. The Process Diagnostics Method (PDM) [4] focuses on rapid, high-level analysis of business processes without requiring domain-specific knowledge. It examines control flow, performance, and organizational aspects based on event logs, providing quick insights. However, PDM lacks explicit iterative refinement of the analysis [23], and is less suitable for long-term process optimization efforts [4].
@@ 2.2 Brief Overview of OCED
(a) OCEL 2.0 metamodel [1]
(b) EKG metamodel [6]
Figure 1: Two metamodels for storing OCED
The OCEL 2.0 metamodel, shown on the left side of Figure 1, defines an event as an atomic action that occurs in a business process, representing the execution of an event type (a.k.a, an activity), such as “place order” in Order-to-Cash process. Each event type can have multiple event attributes, and each event can store different event attribute values. OCEL records objects as separate entities, where each object belongs to an object type. For example, the “order” can be an object type with multiple instances representing different orders. Each object type can have multiple object attributes, and each object can store different values for these attributes at specific points in time. Objects can also be related to other objects through object-to-object relations; for example, an order can contain multiple items. As illustrated in the metamodel, events can be associated with multiple objects and vice versa through event-to-object relations.
@@ 2.3 OCED Extraction Initiatives
Beyond SAP-specific solutions, Xiong et al. introduced a method for extracting OCEL from relational databases using a Virtual Knowledge Graph (VKG) approach [24]. Their method extended the OnProm framework to support both OCEL and the traditional XES format. By leveraging ontology-based data access (OBDA) and the VKG system Ontop, they defined domain ontologies and mappings, enabling event log extraction via SPARQL queries. Their approach was validated using the Dolibarr ERP system, demonstrating how relational data can be transformed into structured OCELs.
@@ 3 The OCPM2 Methodology
This section introduces OCPM2, an Object-Centric Process Mining Methodology defined as an extension of PM2 [23] to facilitate the extraction of Object-Centric Event Data (OCED) [7]. We define our method by extending PM2 to ensure the reuse of a well-established and previously tested methodology for event log extraction, originally designed for traditional process mining. Our approach retains the clarity and efficiency of PM2 while making it suitable for extracting object-centric event data from diverse information systems. We also presented our methodology to 12 domain experts in a workshop, which helped us further confirm the OCPM2 validity.
The methodology consists of twelve stages, where the related stages are grouped into phases, as indicated by dashed rectangles in Figure 2. The process begins with the Planning stage, followed by the Domain Modeling, Log Extraction, and Analysis Iteration phases, and concludes with the Process Improvement & Support stage. Each stage produces specific artifacts, as outlined in the figure. The following sections elaborate on these stages, the generated artifacts, and the defined phases. The filled blue colored phases and stages are identical to those in the PM2 methodology briefly described in Section 2 (we refer readers to the PM2 [23] for further details).
Figure 2: An overview of the OCPM2 methodology extending the PM2 methodology [23]
@@ 3.2 Domain Modeling
The Domain Modeling phase consists of five stages (stages 2 to 6 in Figure 2), identifying and conceptualizing the information needed for the extraction of OCED. As OCED involves multiple object types, a domain model can act as a foundation to facilitate identifying the involved object types, their relations, activities, and the relations between object types and events. This information can be gathered through interviews with business experts, reviewing system documentation, and analyzing the underlying database schema. Without this foundation, the extraction process may result in an incomplete OCED.
@@
The phase begins with Object Type Identification to determine all object types needed to address the questions elicited in the Planing. Because an object type may be required for multiple questions, we propose documenting this using a Question-to-Object Type Matrix (Q2OT), inspired from the bus matrix which is a widely used technqiue in identifying dimensions and related processes in data warehousing [16]. This matrix represents questions as rows and object types as columns, marking intersection cells to indicate the required object types for each question.
@@
Next, in the Conceptual Modeling stage, process analysts and business experts collaborate to define a model ( CM) that captures the relationships between object types. Business experts provide domain knowledge, while process analysts structure this information using a conceptual modeling language to represent key relationships. Among these, recognizing the “is-a” relationship—where one object type is a specialized version of another (e.g., Student is a User)— enables capturing OLAP operations such as drill-down and roll-up in OCPM based on OCEL [17, 13].
@@
The Activity Identification stage follows, eliciting different activities (or event types) to which the identified object types and questions are related. These activities can be identified in various ways, such as analyzing activities directly related to the formulated questions, reviewing system documentation, and consulting business experts.
@@
The Extraction Matrix Definition stage defines the extraction matrix (X), where the relationships between object types and activities are documented. When an “is-a” relationship exists among object types, separate columns can specify whether an activity relates to the supertype class or a subtype class. Various documentation approaches can be adapted based on specific perspectives. From a system perspective, CRUD (Create, Read, Update, and Delete) operations are suitable for specifying the type of relationship between an activity and an object type. However, from a process perspective, CRUD may not fully capture the dynamic nature of object interactions. For example, when a user views a module such as File, CRUD cannot document the role of user properly. Instead, documenting the quantity range of objects for each object type provides a more process-oriented approach. This method captures the expected number of objects per event and supports data validation by ensuring that the extracted log adheres to these expectations. The details of these adaptations can be delegated to design and development teams.
@@
Finally, in the Extraction Prioritization stage, business, process, and system experts collaborate to prioritize the extraction of objects and events based on their i) feasibility, ii) data availability, and iii) importance. Given the potentially large number of object types and activities, extracting all data at once may not be feasible due to data being fragmented across multiple sources and the need to check data quality step-by-step to catch errors early. To manage this, we propose a prioritization artifact as the output of this stage, including a set of prioritized objects and activities ( Pr). Note that new object types can be identified during this phase, requiring the phase to be repeated to have a more holistic understanding of the domain.
@@ 3.3 Log Extraction
The Log Extraction phase in OCPM2 consists of three stages (Stages 7 to 9 in Figure 2) as follows:
In the Implementation stage, developers build pipelines to extract data from source systems, transform them into the desired OCED format, and load them into destination storage. If OCEL is the target format, objects should be processed first based on their list in the extraction matrix. When an “is-a” relationship exists among object types, we propose storing objects at the supertype level with an attribute distinguishing subtypes. This structure follows the single-table inheritance technique used in Object-Relational Mapping (ORM) [8]. This ensures referential integrity and facilitates efficient querying and aggregation, allowing drill-down operations [13, 14]. Subsequently, developers implement extraction logic for object-to-object relationships, determined by the conceptual model (stage 3), in parallel with implementing extraction logic for events based on the extraction matrix. Once event extraction is complete, they implement extraction logic for event-to-object relationships. Please note that similar mapping can be done for EKG [6] and tEKG [11] as the OCED formats are transformable to each other.
@@
Although the log extraction process has been adapted to extract OCED, the remaining stages remain identical to PM2 [23] since the types of applicable process mining techniques are not based on analyzing case-centric or object-centric event logs. However, these phases may include new process mining techniques in the future as the area develops further. Additionally, one can reduce the dimensionality of OCED by converting the log to the traditional format by correlating events to only one object type, a.k.a. flattening, to enable the use of traditional process discovery techniques. The flattening can also help filter the log by identifying similar object type clusters in OCEL [10], meaning that process mining techniques can be applied to a range of case-based and object-centric event logs.
@@ 3.4 Analysis Iteration
The Analysis Iteration phase in PM2 supports gaining insights from extracted event data through two stages: Mining & Analysis and Evaluation. The Mining & Analysis stage includes process model discovery ( PM) from the input log ( L), conformance checking, and process enhancement based on analytical insights ( A) resulting from various analytical techniques, such as performance, and predictive analysis [21]. This can yield compliance findings ( C) and performance findings ( P). The Evaluation stage results in process improvement ideas ( I) or refined/new process-related questions ( ?+), which may prompt further analysis or new data extraction.
@@ 4 Case Study
This section presents the application of the OCPM2 methodology in a process mining project conducted at the Department of Computer and Systems Sciences, Stockholm University. We elaborate on each stage and phase below. In addition, the second and last authors conducted a separate feasibility study demonstrating the applicability of this methodology in a real-world case from the insurance domain [15].
@@ 4.1 Planning
Although we identified numerous questions to guide the project, we present only a subset here to illustrate the application of the methodology. The information systems identified at this stage included 1) the learning management system, Moodle [18], and 2) the platform that records students’ grades.
@@ 4.2 Domain Modeling
Conceptual Modeling: Figure 3 presents the conceptual model using UML notation, a de facto modeling standard widely used in the industry, illustrating the relationships between object types. Nonetheless, the methodology is modeling-language agnostic, and equivalent results can be achieved with alternative notations. This model was developed following the conceptual modeling process described in Section 3.2, based on stakeholder input, system documentation, and database schema analysis. The model includes additional object types beyond those represented in the Q2OT matrix, reflecting the iterative nature of our approach. For example, the object type Teacher was later identified by reviewing the Moodle documentation as a relevant entity within the process. Additionally, User was introduced as a supertype encompassing both Teacher and Student as explained in Section 2.
@@ 4.3 Log Extraction
Implementation: We set up development and test environments to build and validate pipelines using Python that extracted relevant data objects and events, according to the extraction matrix, from Moodle’s relational database. We then transformed the extracted data into OCEL format. When transforming data into OCEL, we defined User as a single object type, with Teacher and Student modeled as subclasses. This approach facilitated the discovery of process models at a general level while enabling drill-down analyses to distinguish teacher- and student-specific process behaviors.
@@
Verification: We developed a script to systematically verify the accuracy of the extracted event data by mapping activities to object types. The resulting dataset contained 8 object types and 607 objects, linked through 7 event types and a total of 20,100 events, as shown in Figure 4. Since the data stems from a single course, all events are associated with this course, meaning the course column aggregates the total number of events.
Figure 4 presents this structured validation process. In this figure, the expected relationship between Group and set assignment grade is missing. This discrepancy arises because Moodle logs a separate set assignment grade event for each student, even when a teacher grades a group assignment collectively. Automated verification helped identify such inconsistencies, highlighting areas for improvement in future log extractions by developers. Furthermore, the relationship between Teacher and Student is not directly visible because the types of objects were recorded at the User level. However, drill-down analysis can reveal these relationships as needed.
@@
In this paper, we have presented only a subset of the case study conducted to demonstrate the application of our proposed methodology. Our full study included additional questions, object types, and events, during which we encountered further challenges in the extraction of OCEL, including the lack of recorded historical temporal data and instances of incomplete data capture.
An example of such challenges is the lack of recording historical temporal data in Moodle for module sections. Moodle only captures the current section for each module, so if a module like a File has been moved, the previous section will not be recorded in the extracted OCEL. We excluded presenting the section in this paper because it was outside the scope of the questions we selected for demonstration.
Another challenge related to incomplete data capture occurs when modules, like File, are imported from one course into another. In these cases, Moodle does not accurately log the importing user’s details, resulting in the user information for these imported modules being unspecified. This incomplete logging limits the completeness of the extracted OCEL.
@@ 4.4 Analysis Iteration
For Q1, Figure 5 illustrates an Object-Centric Directly-Follows Graph (OC-DFG), discovered using the PM4Py library. This graph was generated by filtering the extracted log, drilling down page objects into individual page names, and unfolding view page events for each page using the processmining library [14]. The graph visualizes the sequential flow of actions during the course from the perspectives of the User and Course object types.
@@
In this process, the page names for given codes are as follows: A1P1: History of BPM, A1P2: Business Process Models, A1P3: Business Process Enactment, A1P4: Syntax and Semantics, A1P5: Control-flow Patterns, A1P6: Business Process Complexity, and L1P1: Introduction to Process Tree. The model reveals that students sequentially accessed five pages: A1P1 to A1P5. Additionally, a loop between A1P5, A1P6, and L1P1 suggests students frequently revisit these pages, indicating potential room for improvement.
@@ 5 Conclusion
To validate our approach, we applied OCPM2 in an educational setting, extracting and analyzing OCELs from a learning management system and an administrative grading system, which cannot be shared due to privacy concerns. Our results demonstrated the practical usefulness of OCPM2, i.e., its ability to effectively reveal student navigation patterns and assignment submission processes.
The case study exposed limitations in system logs, particularly in capturing accurate object-to-activity relationships. A key issue was the missing link between Group and set assignment grade, causing Moodle to log separate grading events per student, even for collectively graded assignments. This discrepancy resulted in a mismatch between the number of students who submitted the assignment and those involved in the grading process. Automated verification identified these inconsistencies, helping developers extract correct event logs in future extractions. Future research will focus on enhancing automated verification mechanisms, refining multi-dimensional process analysis operations, and expanding OCPM2’s applicability across other domains, such as healthcare and financial services. We also plan to develop a tool to support OCED extraction based on OCPM2. This tool will enable organizations to document and manage OCED extraction for different projects.